A 52.53 gram sample of iron (with a heat capacity of 0.450 J/gºC) is heated to 100.0 ºC. It is then transferred to a coffee cup calorimeter containing 49.57 g of water (specific heat of 4.184 J/ gºC) initially at 20.63 ºC. If the final temperature of the system is 23.59ºC, how much heat was absorbed by the calorimeter? (Please give the absolute value)
Answer #1

Solution :-

Lets calculate the heat loss by the iron

Mass of iron = 52.53 g

Specific heat of iron = 0.450 J/gC

Change in temperature = 23.59 C – 100 C = -76.41 C

q iron =-m*c*delta T

q=-52.53 g * 0.450 J per g C * (-76.41 C)

q=1806.22 J

Now we can find the heat absorbed by the water

q water = m*c*delta T

               = 49.57 g* 4.184 J per g C * (23.59 C-20.63 C)

               = 613.9 J

Heat lost by iron = heat gain by water + heat gain by calorimeter

Heat gain by calorimeter = heat lost by iron – heat gain by water

                                             = 1806.22 J -613.9 J

                                             = 1192 J

Therefore heat absorbed by the calorimeter is 1192 J